OpenStudy (anonymous):

Determine the solution to the given equation. 4+3y=6y-5

OpenStudy (anonymous):

4+3y=6y-5 4+3y+5=6y-5+5 9+3y=6y 9+3y-3y=6y-3y 9=3y 3y = 9 3y/3 = 9/3 y = 3